# Correcting Nonresponse Bias Using Panel Data

### v3

@@ 3.4 Oaxaca-Blinder-Kitagawa Decomposition
$\hat{\Delta}=\sum_{i=N_{w}+1}^{N}\frac{\Phi(X_{i}\hat{\beta}_{m})}{N_{m}}-\sum_{i=1}^{N_{w}}\frac{\Phi(X_{i}\hat{\beta}_{w})}{N_{w}},$ (12)
where our estimates of $(\hat{\beta}_{w},\hat{\beta}_{m})$ differ by estimation method. The estimated gender-intention gap from our preferred specification is 18.3 percentage points, with a naive probit producing an estimate of 20.2 percentage points. Differences between the corrected and uncorrected model are due to the inclusion of the $\rho$ parameter in the corrected model (with $\rho=0$ fixed in the uncorrected model) which is positive if early respondents have higher intention and negative if they have lower intention. Our estimates of $\rho$ in the current specification are insignificant, suggesting limited nonresponse bias, though we see the same pattern with the signs of our point estimates here as we did in the no-controls specification shown in Figure 1. Along these same lines, our richer specification repeats the finding in Figure 1 that the uncorrected gap point estimate is larger than the corrected estimate, but here we fail to reject that they are the same in the population (p=0.449).
The difference in corrected estimates in Section 3.3 and the corrected estimates in Table 2 is due to the inclusion of controls and observations for requests three through seven, which were observed in some years but not others. The richer nonresponse correction specification in this section approximates a nonparametric specification that would fit a curve through local responses as in Figure 1 for each value of $X_{i}$, where the slope of the curve is determined by a gender-specific $\rho$ that is held fixed across $X_{i}$. For both men and women, $\rho$ is attenuated in the richer model, suggesting that observed covariates account for both late responses and differences in intention between late and early responses. Nonetheless, the same general pattern from Figure 1 persists, with positive selection for men ($\hat{\rho}_{m}>0)$ and negative selection for women $(\hat{\rho}_{f}<0)$, though these estimates are not statistically significant in this specification.
@@ 4 Conclusion
We illustrate our method by estimating the gender gap in entrepreneurial intentions among undergraduate students at the University of Wisconsin-Madison. To do so, we construct a panel dataset that contains a row for each student for each data request they receive, with their accumulated requests recorded in all rows and their responses recorded only in rows representing time periods after they respond. We find a large gender gap both without correcting for nonresponse and when correcting for nonresponse using requests as instruments in a van1981demand model, with the uncorrected estimates strongly rejecting the point estimate of the average gender gap obtained when correcting for nonresponse. In an additional application in Appendix B, we accurately and precisely estimate ground truth averages from administrative data for five out of six labor market indicators using only aggregated statistics from early and late responses to the 45% response-rate Norway in Corona Times survey. Our empirical results suggest that correcting for nonresponse using nonrandom request instruments in subject contact panel data can improve inferences, while posing relatively small challenges for data collection and analysis.

### v4

@@ 3.4 Oaxaca-Blinder-Kitagawa Decomposition
$\hat{\Delta}=\sum_{i=N_{w}+1}^{N}\frac{\Phi(X_{is}\hat{\beta}_{m})}{N_{m}}-\sum_{i=1}^{N_{w}}\frac{\Phi(X_{is}\hat{\beta}_{w})}{N_{w}},$ (12)
where our estimates of $(\hat{\beta}_{w},\hat{\beta}_{m})$ differ by estimation method. The estimated gender-intention gap from our preferred specification is 18.3 percentage points, with a naive probit producing an estimate of 20.2 percentage points. Differences between the corrected and uncorrected model are due to the use of only observations with $t=T$ for the naive probit and the inclusion of the $\rho$ parameter in the corrected model (with $\rho=0$ fixed in the uncorrected model) which is positive if early respondents have higher intention and negative if they have lower intention. Our estimates of $\rho$ in the current specification are insignificant, suggesting limited nonresponse bias, though we see the same pattern with the signs of our point estimates here as we did in the no-controls specification shown in Figure 1. Along these same lines, our richer specification repeats the finding in Figure 1 that the uncorrected gap point estimate is larger than the corrected estimate, but here we fail to reject that they are the same in the population (p=0.449).
The difference in corrected estimates in Section 3.3 and the corrected estimates in Table 2 is due to the inclusion of controls and observations for requests three through seven, which were observed in some years but not others. The richer nonresponse correction specification in this section approximates a nonparametric specification that would fit a curve through local responses as in Figure 1 for each value of $X_{is}$, where the slope of the curve is determined by a gender-specific $\rho$ that is held fixed across $X_{i}$. For both men and women, $\rho$ is attenuated in the richer model, suggesting that observed covariates account for both late responses and differences in intention between late and early responses. Nonetheless, the same general pattern from Figure 1 persists, with positive selection for men ($\hat{\rho}_{m}>0)$ and negative selection for women $(\hat{\rho}_{w}<0)$, though these estimates are not statistically significant in this specification.
@@ 4 Conclusion
We illustrate our method by estimating the gender gap in entrepreneurial intentions among undergraduate students at the University of Wisconsin-Madison. To do so, we construct a panel dataset that contains a row for each student for each data request they receive, with their accumulated requests recorded in all rows and their responses recorded only in rows representing time periods after they respond. We find a large gender gap uncorrected estimates strongly rejecting the point estimate of the average gender gap obtained when correcting for nonresponse. In an additional application in Supplemental Appendix B, we accurately and precisely estimate ground truth averages from administrative data for five out of six labor market indicators using only aggregated statistics from early and late responses to the 45% response-rate Norway in Corona Times survey. Our empirical results suggest that correcting for nonresponse using nonrandom request instruments in subject contact panel data can improve inferences, while posing relatively small challenges for data collection and analysis.